OpenStudy (anonymous):

If the first step in the solution of the equation 2x - 8 = 5x + 3 is "subtract 2x," then what should the next step be?

HanAkoSolo (jamierox4ev3r):

after you subtract 2x from both sides subtract 3 from both sides

HanAkoSolo (jamierox4ev3r):

then you are left with -11=3x

OpenStudy (anonymous):

would i subtract 2x from 5x?

OpenStudy (anonymous):

There isn't really one right answer to this, as you can approach it from several different ways. I would subtract 3 from both sides, I think thats the simplest way, as you'll be one step away from isolating x.

HanAkoSolo (jamierox4ev3r):

yes @Zacharymartin

HanAkoSolo (jamierox4ev3r):

then, your final answer is \[-11\over 3\]

HanAkoSolo (jamierox4ev3r):

does that make sense @Zacharymartin ?

OpenStudy (anonymous):

so when doing these kinda problems with a varible on both sides would i subtract from the other variable after?

OpenStudy (anonymous):

2x - 8 = 5x + 3 -2x -2x ---- ----- -8 = 3x + 3 -3 -3 --- ------ -11 = 3x ---- ---- 3 3 \[x = -\frac{ 11 }{ 3 }\]

OpenStudy (anonymous):

2x - 8 = 5x + 3 (subtract 2x) -8 = 5x - 2x + 3 (subtract 3) -8 - 3 =5x - 2x -11 = 3x -11/3 = x

OpenStudy (anonymous):

so i need the 2 varibles on one side?